OpenStudy (anonymous):

The square of a number decreased by 25 is -9. A. Find the x-intercepts. B. Find the vertex of the graph. C. What shape is the graph? Describe what you see

OpenStudy (acxbox22):

x^2-25=-9 x^2-16=0 (x-4)(x+4)=0 x int: (4,0) and (-4,0)

OpenStudy (anonymous):

Thank you.